For a standardized psychology examination intended for psychology majors, the historical data show that scores have a mean of 500 and a standard deviation of 170. The grading process of this year's exam has just begun. The average score of the 40 exams graded so far is 490. What is the probability that a sample of 50 exams will have the mean score of 490 or more if the exam scores follow the same distribution as in the past?
Carry your intermediate computations to at least four decimal places. Round your answer to at least three decimal places.

Answer by ewatrrr(24785)


Hi
 z =   (x̄ - μ) /(170/√50) = -10/(170/√50) = -.4159
What is the probability that a sample of 50 exams will have the mean score
of 490 or more if the exam scores follow the same distribution as in the past?
Continous curve: P( z ≤ -.4159) = .3387
P( z ≥ -.4159) =  1 -  .3387 = .6613
Wish You the Best in your Studies.